OpenStudy (anonymous):

Solve. 6 – 3x = 4(3 – x) – 6 A. 6 B. –6 C. 0 D. all real numbers

undeadknight26 (undeadknight26):

wrong x = 0...

OpenStudy (anonymous):

6-3x=12-4x-6 got dis step?

OpenStudy (anonymous):

now we have to bring x terms to one side nd constant no.s should be moved to the other side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6–3x=4(3–x)–6 lets first distribute 4 over the parantheses so 6-3x=12-4x-6 now simplify the right side 6-3x=6-4x now add 4x to both sides to move the variables to the left 6+x=6 substract 6 from both sides x=0

undeadknight26 (undeadknight26):

Give medal to @insanitycat she did the real work and showed you how to do it... thx!!!